M=1/2
R(0,0)
Write in slope intercept form

Answer by Thanatos(1)

This problem can be solved by using this equation...
y-(y1)=m(x-(x1))
where (y1) and (x1) is a point located on your graph or information given. Eg, a point like (3,9).. would have a (x1) point of 3 and a (y1) point of 9.
And also m is your gradient.
Here you have a point equal to (0,0) so (x1) is 0 and (y1) is 0.
m, which is also given to you is 1/2.
So plugging these into the equation about and you get...
y-0=1/2(x-0) (which simplifies as)
y=(1/2)x (which simplifies as)
y=x/2